OpenStudy (xmoses1):

what is 2 times root 3 squared? is it six?

OpenStudy (andrewkaiser333):

square root of 3 is 3

OpenStudy (andrewkaiser333):

times 2

OpenStudy (andrewkaiser333):

equals 6

OpenStudy (johnweldon1993):

\[(2 * \sqrt{3})^2\] or is it \[2 * \sqrt{3}^2\] ?

OpenStudy (andrewkaiser333):

ah i am doing it wrong

OpenStudy (xmoses1):

it is the second! sorry

OpenStudy (andrewkaiser333):

square root of 9 is 3

OpenStudy (andrewkaiser333):

times 2 equals 6

OpenStudy (johnweldon1993):

Yes okay so then it is the second The SQUARE ROOT of 3....raised to the second power.....eliminates the square root. √3^2 = 3 So you just have 2 * 3 which yes...is 6

OpenStudy (andrewkaiser333):

so same answer my bad

OpenStudy (andrewkaiser333):

yes it is 6

OpenStudy (xmoses1):

Alright that is what i thought i just wanted to ask someone who knew for sure:) Thank you!

OpenStudy (andrewkaiser333):

np

OpenStudy (andrewkaiser333):

you can give best response if you want i do not really care if you do i am just trying to help

OpenStudy (xmoses1):

I will make sure that you get one:)

OpenStudy (andrewkaiser333):

ok

OpenStudy (andrewkaiser333):

gttg bye

OpenStudy (andrewkaiser333):

gtg*